# Preference for Verifiability

### v1

@@ 7 Relation to Literature
Our model is a special case of the classic ambiguity models of maxmin expected utility [32], and Choquet expected utility [39]. Most closely related to the present paper is the dual self model of ambiguity [24]. In this model, decision makers maximize an objective $U(a)=\max_{P\in\mathscr{P}}\min_{p\in P}\int u\circ adp$ where $\mathscr{P}$ is a set of sets of priors. The interpretation of the model is that there are two selves, a pessimistic one and an optimistic one. The pessimistic one evaluates the acts according to the maxmin expected utility model [32] by choosing the worst possible prior. However, the optimistic self determines the set of available priors that the pessimist can choose from. The present model is a special case of the dual self model. In a dual self interpretation of our model, the maximizer chooses a minimal cover of verifiable events. Each such cover corresponds to a set of priors via the restriction that a prior $p$ needs to fulfill $p(E)=\mu(E)$ for all events $E$ in the cover. In other words, to evaluate an act, the minimizer can choose to assign the entire probability of a verifiable event to the worst state of the verifiable event. Relative to the characterization of [24], our model provides conditions under which the maximization and minimization steps can be written inside the expectation, i.e., we provide a way to determine ambiguity-neutral events.
@@
If the set of verifiable events is exogeneously given, a characterization using cominimum additivity [35, 36] instead of comonotonic independence is possible. Cominimum additivity with respect to a set of events $\mathscr{V}$ requires that all acts that agree on the worst states in all of the events in $\mathscr{V}$ have to fulfill the independence axiom. Cominimum additivity provides intermediate cases between full independence and comonotone additivity.
@@
Preference for verifiability is a form of information preference since the decision maker cares about how much information is ex post available about which consequence has been achieved. Information preferences are often modeled using a two stage approach [37, 40, 27]. Given our decision model, the first stage would correspond to the verifiable events and the second stage to the final consequences and only the first stage is observed. If verifiable events $\mathscr{V}$ are objectively given and both ex-ante and ex-post probability distributions over consequences are objective, it is possible to perform our analysis using preferences over $\Delta\mathscr{V}\times\Delta\mathscr{X}^{\mathscr{V}}$. However, this way of modeling would severely restrict the explanatory power of our model. It does not allow to infer the verifiable events from behavior but requires the analyst to have data on the ex-post beliefs of the decision maker. In the context of our example, this would defeat the main purpose of the paper since ex-post the stakeholder and decision maker would agree on the probability distribution of the CO2 emissions.
@@
The present paper also provides a decision theoretic foundation for the standard notion of verifiability employed in contract theory starting with [23]: verifiable events are closed under intersection but not necessarily under relative complements. Our application shows that this notion can be productively used to model greenwashing.
@@
Our research also relates to definitions [26] and formal models of greenwashing [42] and of green products [34]. Unlike previous modeling attempts, the present paper provides a purely behavioral definition of greenwashing versus verification-seeking behavior. We do not require unobservable model components such as incorrect consumer beliefs due to deception or detailed information about the interaction between consumers and firms. This yields a very parsimonious model of greenwashing which does not require any information about the behavior of consumers of greenwashed products – preference data of the firm over “green” policies fully identifies the model.
@@ 8 Discussion
When consequences are not directly observable to decision makers, it is plausible that these decision makers deviate from expected utility. The present paper provides a starting point for the analysis of such deviations. Specifically, decision makers may for example care about verifying or obfuscating what consequence has been achieved in the presence of verifiable information. As our axiomatic analysis shows, unobservable consequences may provide a rationalization for an extreme form of ambiguity preferences. Crucially, such behavior can be identified from preferences over acts alone and it is not necessary to introduce additional primitives. In particular, we neither need to enrich the outcome space by ex-post beliefs over consequences nor do we need to impose any information structure of what is verifiable.
Starting with [41], there have been attempts to increasingly rely on CCR to achieve CO2 emission reductions. These attempts have been criticised as “inappropriately verified” [[], p.6]new_corporate_2023. Our comparative statics results suggest that unless perfect ex-post verifiability is implemented, policies aimed at increasing verifiability of consequences may not be suitable to reduce the welfare losses from verification-seeking and obfuscation-seeking behavior. In such cases, traditional environmental policies that directly incentivize expected CO2 emission reductions could be more suitable than a combination of CCR and transparency policies.
Appendix A Proof of Theorem 1
@@
By comonotonic independence and the biseparable utility representation we obtain [[, see Proposition 2 of]]ghirardato_subjective_2003 a representation $U(a)=\int(u\circ a)d\mu$ where $\mu$ is a capacity and the integral is in the sense of Choquet. Let $m^{\mu}$ be the Möbius inverse of $\mu$, i.e., $m^{\mu}(E)=\sum_{A\subseteq E}(-1)^{|A|-1}\mu(A)$. The Choquet integral expressed in terms of the Möbius inverse [[, see ], p. 235]grabisch_set_2016 is given by
We now prove the key lemma which relates properties of $m^{\mu}$ to whether a set is critical.
@@
  • if $m(S)>0$ then $S$ is critical and there does not exist a cover of critical sets that are all strict subsets of $S$.
@@
Case $n=1$: Since $\mu$ is a capacity, $m(S)+m(\emptyset)=\mu(S)\geq\mu(\emptyset)=m(\emptyset)=0$. Thus $m(S)\geq 0$. If $m(S)=\mu(S)>0=\mu(\emptyset)$, then since $S$ contains a single element, $S$ is critical and has no critical subsets.
Case $n>1:$ Suppose for all sets $S$ of size n1n-1 or smaller the induction hypothesis holds. We distinguish the case that $m(S)<0$ from the case that $m(S)>0$ and there exists a cover of $S$ of critical events that are strict subsets of $S$ and derive a contradiction for each case.
@@
    We first show that $S$ is critical and there exists a cover of critical subsets of $S$. For every $s\in S$ we have by monotonicity of the capacity that $\mu(S)\geq\mu(S-\{s\})$. It follows from the definition of $m$ that $m(S)+\mu(S-\{s\})+\sum_{E\subset S:s\in E}m(E)=\mu(S)\geq\mu(S-\{s\})$ and thus $m(S)+\sum_{E\subset S:s\in E}m(E)\geq 0$. Therefore for some $E\subset S$ containing $s$ we have that $m(E)>0$ and by the induction hypothesis $E$ is critical. It follows that every $s\in S$ is contained in a critical event and since by critical event modularity critical events are closed under unions, $S$ is also critical.
    Since $n\geq 2$ and there exists a cover of critical subsets of $S$ and critical subsets are closed under unions, we can find a cover $\{A,B\}$ of two critical subsets of $S$ with ABA-B and BAB-A nonempty. Then by critical event modularity, $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$. Thus, by the definition of $m$, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$. This is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}m(E)=0$. Because by assumption $m(S)<0$, this is only possible if $m(E)>0$ for some $E\subset S$, $E\not\subseteq A,B$. But from critical event modularity follows that $\mu(E)+\mu(E\cap A\cap B)=\mu(A\cap E)+\mu(B\cap E)$ and thus $\sum_{F\subset E:F\not\subseteq A,F\not\subseteq B}m(F)<0$, contradicting the induction hypothesis.
  • Suppose for sake of contradiction that $m(S)>0$ and there exists a cover of critical subsets of $S$. Then $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$ by critical event modularity. Thus, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$ which is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}=0$. But this is only possible if $m(E)<0$ for some $E\subset S$, contradicting the induction hypothesis.
@@
By Lemma 1, $E$ and $F$ are critical. It follows from critical event modularity that $E\cap F$ is critical. Suppose $m(G)=0$ for all $G\subseteq E\cap F$ such that $s\in G$. Then $\mu(E\cap F)=\sum_{G\subseteq E\cap F}m^{\mu}(G)=0+\sum_{G\subseteq(E\cap F)-\{s\}}=\mu((E\cap F)-\{s\})$, contradicting that $E\cap F$ is critical. Thus, there exists some $G\subseteq E\cap F$ such that $m^{\mu}(G)>0$. ∎
@@
Appendix B Proof of Proposition 1
@@
From the uniqueness properties of the biseparable preferences follows that $U^{1}=\theta U^{2}+\phi$ and $u^{1}=\theta u^{2}+\phi$. What is left to show is the relation between the verifiable sets $\mathscr{V}^{1}$ and $\mathscr{V}^{2}$. Note that the critical events in the representation are exactly the sets $E$ for which $m^{\mu}(E)>0$ or for which there exists a cover of sets $E_{1},\ldots,E_{n}\subseteq E$ such that $\forall i:m^{\mu}(E_{i})>0$.
$\mathscr{V}^{1}\subseteq\mathscr{V}^{2}$: If $DV=cl_{\cup}(\mathscr{V}^{1})-cl_{\cup}(\mathscr{V}^{2})$ is nonempty, then for some $\in DV, VV$ is critical in $U^{1}$. We show that $V$ cannot be critical in $U^{2}$: If $V$ is critical in $U^{2}$, then either $m^{\mu^{2}}(V)>0$ and thus $V\in\mathscr{V}^{2}$ or there exists a cover $V_{1},\ldots,V_{n}$ such that $\forall i:m^{\mu^{2}}(V_{i})>0$. If the latter is the case, then $V_{1},\ldots,V_{n}\in\mathscr{V}^{2}$ and thus $V\in cl_{\cup}(\mathscr{V}^{2})$. It follows that $V\in DV$ cannot be critical in $U^{2}$. But if $V$ is critical in $U^{1}$ but not in $U^{2}$, then the preferences cannot be identical.
$\mathscr{V}^{2}\subseteq\mathscr{V}^{1}$ follows by a symmetric argument.
@@
Appendix C Proof of Theorem 2

### v2

@@ 7 Relation to Literature
Our model is a special case of Choquet expected utility [undefap]. While CEU has lost some of its popularity as a model of ex-ante ambiguity aversion, our interpretation as a preference for verifiability provides a clear normative motivation for it in the context of ex-post uncertainty and identifies a choice domain in which it is the natural representation. In particular, as discussed in the axioms section, comonotonic independence seems a natural starting point for other, less extreme versions of verifiability preferences than ours.
@@
If the set of verifiable events is exogenously given, a characterization using cominimum additivity [undefak, undefal] instead of comonotonic independence is also possible. Cominimum additivity with respect to a set of events $\mathscr{V}$ requires that all acts that agree on the worst states in all of the events in $\mathscr{V}$ have to fulfill the independence axiom. Cominimum additivity provides intermediate cases between full independence and comonotone additivity.
Closely related to the representation of the present paper is the dual self model of ambiguity [undefy]. In this model, decision makers maximize an objective $U(a)=\max_{P\in\mathscr{P}}\min_{p\in P}\int u\circ adp$ where $\mathscr{P}$ is a set of sets of priors. The interpretation of the model is that there are two selves, a pessimistic one and an optimistic one. The pessimistic one evaluates the acts according to the maxmin expected utility model [undefag] by choosing the worst possible prior. However, the optimistic self determines the set of available priors that the pessimist can choose from. Our two main models are special cases of the dual self model. In a dual self interpretation of our model, the maximizer chooses a minimal cover of verifiable events. Each such cover corresponds to a set of priors via the restriction that a prior $p$ needs to fulfill $p(E)=\mu(E)$ for all events $E$ in the cover. In other words, to evaluate an act, the minimizer can choose to assign the entire probability of a verifiable event to the worst state of the verifiable event. Relative to the characterization of [undefy], our model provides conditions under which the maximization and minimization steps can be written inside the expectation.
@@
Preference for verifiability is a form of information preference since the decision maker cares about how much information is ex post available about which consequence has been achieved. Information preferences are often modeled using a two stage approach [undefam, undefaq, undefab]. Given our decision model, the first stage would correspond to the verifiable events and the second stage to the final consequences and only the first stage is observed. If verifiable events $\mathscr{V}$ are objectively given and both ex-ante and ex-post probability distributions over consequences are objective, it is possible to perform our analysis using preferences over $\Delta\mathscr{V}\times\Delta\mathscr{X}^{\mathscr{V}}$. However, this way of modeling would severely restrict the explanatory power of our model. It does not allow to infer the verifiable events from behavior but requires the analyst to have data on the ex-post beliefs of the decision maker. In the context of our example, this would defeat the main purpose of the paper since ex-post the stakeholder and decision maker would agree on the probability distribution of the CO2 emissions.
@@
The present paper also provides a decision theoretic foundation for the standard notion of verifiability employed in contract theory starting with [undefx]: verifiable events are closed under intersection but not necessarily under relative complements. Our application shows that this notion can be productively used to model greenwashing.
@@
Our research also relates to definitions [undefaa] and formal models of greenwashing [undefas] and of green products [undefai]. Unlike previous modeling attempts, the present paper provides a purely behavioral definition of greenwashing versus certification-seeking behavior. We do not require unobservable model components such as incorrect consumer beliefs due to deception or detailed information about the interaction between consumers and firms. This yields a very parsimonious model of greenwashing which does not require any information about the behavior of consumers of greenwashed products – preference data of the firm over “green” policies fully identifies the model.
@@ 8 Discussion
When consequences are not directly observable to decision makers, it is plausible that these decision makers deviate from expected utility. We provide a starting point for the analysis of such deviations. As our axiomatic analysis shows, unobservable consequences can rationalize a form of ambiguity preferences. Specifically, seemingly ambiguity averse behavior can be induced from the desire to be able to prove that good consequences have been reached and seemingly ambiguity loving behavior can be induced from obfuscation seeking behavior. These models only provide extreme starting points for a more general analysis of ex-post verifiability within the CEU framework.
Our reinterpretation of CEU opens a wide field of applications for CEU when verifiability of consequences matters, such as in environmental economics. Starting with [undefar], there have been attempts to increasingly rely on CCR to achieve CO2 emission reductions. These attempts have been criticised as “inappropriately verified” [[], p.6]new_corporate_2023. Our comparative statics results suggest that —unless perfect ex-post verifiability is implemented— even in our simplistic model application the welfare effect of policies aimed at increasing verifiability of consequences are indeterminate. In such cases, traditional environmental policies (taxation of fossil fuels, energy efficiency mandates, etc.) that directly incentivize actions rather than consequences avoid this indeterminacy altogether, since their welfare effects do not depend on the structure of verifiable events.
Acknowledgements
The author gratefully acknowledges support from the Japan Society for the Promotion of Science (JSPS) under grant number 25K05000.
Appendix A Proof of Corollary 1
We prove sufficiency. By comonotonic independence and the biseparable utility representation we obtain [[, see Proposition 2 of]]ghirardato_subjective_2003 a representation $U(a)=\int(u\circ a)d\mu$ where $\mu$ is a capacity and the integral is in the sense of Choquet.
Let $m^{\mu}$ be the Möbius inverse of $\mu$, i.e., $m^{\mu}(E)=\sum_{A\subseteq E}(-1)^{|A|-1}\mu(A)$. By Proposition 2 of [undefz], the Choquet integral can be expressed in terms of the Möbius inverse [[, for a comprehensive overview, see ], p. 235]grabisch_set_2016:
such that $m^{\mu}(\emptyset)=0$, $\sum_{E\in\mathscr{E}}m^{\mu}(E)=1$, and all partial sums across events contained in an arbitrary event are nonnegative.
[…]
Appendix B Proof of Theorem 1
@@
We prove sufficiency. Via the proof of Corollary 1 we have a representation by a Choquet integral:
$\int(u\circ a)d\mu=\sum_{E\in\mathscr{E}}m^{\mu}(E)\min_{s\in E}u(a(s)).$ (B.9)
We first prove that ambiguity neutrality above sensitive events implies modularity of $\mu$ above sensitive events.
Definition 11 (Modularity above sensitive events).
A capacity $\mu$ is modular above sensitive events if for all sensitive events $E$ and $F$, and all events $A\subseteq E\cup F$:
$\mu(A\cap E\cap F)+\mu(A)=\mu(A\cap E)+\mu(A\cap F).$ (B.10)
Lemma 1 (Modularity).
[…]
Proof.
Without loss of generality, assume $u(\gamma)-u(\beta)=1=u(\gamma^{\prime})-u(\beta^{\prime})$ and $f=\gamma A\cap E\beta\sim\gamma^{\prime}A\cap F\beta^{\prime}=g$. It follows that
$U(1/2f\oplus 1/2g)=$ $1/2\mu(A\cap E\cap F)+1/2\mu(A)+\frac{u(\beta)+u(\beta^{\prime})}{2}$ (B.11)
$U(f)=$ $\mu(A\cap E)+u(\beta)$ (B.12)
$U(g)=$ $\mu(A\cap F)+u(\beta^{\prime})$ (B.13)
are all equal by proof sufficiency. But then $2\cdot U(1/2f\oplus 1/2g)=U(f)+U(g)$ or after substitution:
$\mu(A\cap E\cap F)+\mu(A)+u(\beta)+u(\beta^{\prime})=\mu(A\cap E)+u(\beta)+\mu(A\cap F)+u(\beta^{\prime})$ (B.15)
and thus $\mu$ is indeed modular above $E\cup F$.
From modularity of $\mu$ above $E\cup F$ follows that the sensitive events are not only closed under intersections, but also under unions. This is because modularity implies additivity of the Choquet integral above sensitive events $E$ and $F$: $U(1/2\gamma_{E\cap A}\beta\oplus 1/2\gamma^{\prime}_{F\cap A}\beta^{\prime})=1/2U(\gamma_{E\cap A}\beta)+1/2U(\gamma^{\prime}_{F\cap A}\beta^{\prime})$. ∎
We now prove the key lemma which relates properties of $m^{\mu}$ to whether a set is sensitive.
@@
  • if $m(S)>0$ then $S$ is sensitive and there does not exist a cover of sensitive sets that are all strict subsets of $S$.
@@
Case $n=1$: Since $\mu$ is a capacity, $m(S)+m(\emptyset)=\mu(S)\geq\mu(\emptyset)=m(\emptyset)=0$. Thus $m(S)\geq 0$. If $m(S)=\mu(S)>0=\mu(\emptyset)$, then since $S$ contains a single element, $S$ is sensitive and has no sensitive subsets.
Case $n>1:$ Suppose for all sets $S$ of size n1n-1 or smaller the induction hypothesis holds. We distinguish the case that $m(S)<0$ from the case that $m(S)>0$ and there exists a cover of $S$ of sensitive events that are strict subsets of $S$ and derive a contradiction for each case.
@@
    We first show that $S$ is sensitive and there exists a cover of sensitive subsets of $S$. For every $s\in S$ we have by monotonicity of the capacity that $\mu(S)\geq\mu(S-\{s\})$. It follows from the definition of $m$ that $m(S)+\mu(S-\{s\})+\sum_{E\subset S:s\in E}m(E)=\mu(S)\geq\mu(S-\{s\})$ and thus $m(S)+\sum_{E\subset S:s\in E}m(E)\geq 0$. Therefore for some $E\subset S$ containing $s$ we have that $m(E)>0$ and by the induction hypothesis $E$ is sensitive. It follows that every $s\in S$ is contained in a sensitive event and since by Lemma 1 sensitive events are closed under unions, $S$ is also sensitive.
    Since $n\geq 2$ and there exists a cover of sensitive subsets of $S$ and sensitive subsets are closed under unions, we can find a cover $\{A,B\}$ of two sensitive subsets of $S$ with ABA-B and BAB-A nonempty. Then by modularity above sensitive events, $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$. Thus, by the definition of $m$, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$. This is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}m(E)=0$. Because by assumption $m(S)<0$, this is only possible if $m(E)>0$ for some $E\subset S$, $E\not\subseteq A,B$. But from proof sufficiency follows that $\mu(E)+\mu(E\cap A\cap B)=\mu(A\cap E)+\mu(B\cap E)$ and thus $\sum_{F\subset E:F\not\subseteq A,F\not\subseteq B}m(F)<0$, contradicting the induction hypothesis.
  • Suppose for sake of contradiction that $m(S)>0$ and there exists a cover of sensitive subsets of $S$. Then $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$ by modularity above sensitive events. Thus, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$ which is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}m(E)=0$. But this is only possible if $m(E)<0$ for some $E\subset S$, contradicting the induction hypothesis.
@@
By Lemma 2, $E$ and $F$ are sensitive. It follows from proof conjunction that $E\cap F$ is sensitive. Suppose $m(G)=0$ for all $G\subseteq E\cap F$ such that $s\in G$. Then $\mu(E\cap F)=\sum_{G\subseteq E\cap F}m^{\mu}(G)=0+\sum_{G\subseteq(E\cap F)-\{s\}}m^{\mu}(G)=\mu((E\cap F)-\{s\})$, contradicting that $E\cap F$ is sensitive. Thus, there exists some $G\subseteq E\cap F$ such that $m^{\mu}(G)>0$. ∎
@@
Appendix C Proof of Corollary 2
Proof.
We first prove the following general lemma.
Lemma 5.
[…]
$\mu(E\cup F)+\mu(E\cap F)\geq\mu(E)+\mu(F).$ (C.17)
Proof.
Note that if an event $E$ is sensitive, removing any state from it decreases $\mu(E)$. If it is not sensitive, there exists a state $s$ such that $\mu(E)=\mu(E-\{s\})$.
If $E$ and $F$ are sensitive, supermodularity directly follows from setting $A=E\cup F$ in the definition of modularity above sensitive events.
If one of the two events is not sensitive because a state $s$ can be removed, removing such states from EFE-F and FEF-E does not change the RHS and does not increase the LHS of (C.17) because $\mu$ is a capacity. Therefore it is sufficient to prove supermodularity in the case in which all states that can be removed from non-sensitive events $E$ and/or $F$ are within $E\cap F$.
Suppose $E$ is not sensitive and we could remove state $s\in E\cap F$. Define $E^{\prime}=E-\{s\}$. $\mu(E)=\mu(E^{\prime})$. Thus, again the RHS of (C.17) stays constant if we remove $s$, i.e., $\mu(E)+\mu(F)=\mu(E^{\prime})+\mu(F)$ and the LHS may decrease. Thus, supermodularity with respect to $E$ and $F$ is implied by supermodularity with respect to $E^{\prime}$ and $F$. We can iterate this step until we reach one of two cases: Either $E$ becomes empty and supermodularity holds trivially or $E$ becomes sensitive. If $E$ becomes sensitive, we repeat the above process for event $F$. At the end of this process, either $F$ is empty and $E$ is sensitive, or both $E$ and $F$ are sensitive. In the former case, supermodularity is trivial because the LHS and RHS are identical expressions and in the latter case supermodularity follows from setting $A=E\cup F$ in the definition of modularity above sensitive events. ∎
Since an expected certification utility satisfies modularity above sensitive events (Lemma 1 in the Proof of Theorem 1) and sensitive events are closed under unions and intersections (via proof conjunction and Lemma 1), the capacity $\mu$ is supermodular.
By [undefah], a capacity is supermodular if and only if its associated Choquet integral is superadditive, i.e., for all bounded functions $h_{1},h_{2}$:
$\int(h_{1}+h_{2})\,d\mu\geq\int h_{1}\,d\mu+\int h_{2}\,d\mu.$ (C.18)
Now let $f\sim g$, i.e., $U(f)=U(g)$. The mixed act $\frac{1}{2}f+\frac{1}{2}g$ maps each state $s$ to an outcome such that $u\!\left(\frac{1}{2}f(s)\oplus\frac{1}{2}g(s)\right)=\frac{1}{2}u(f(s))+\frac{1}{2}u(g(s))$. Therefore, by superadditivity of the Choquet integral:
$U\!\left(\tfrac{1}{2}f+\tfrac{1}{2}g\right)$ $=\int\left[\tfrac{1}{2}(u\circ f)+\tfrac{1}{2}(u\circ g)\right]d\mu\geq\tfrac{1}{2}\int(u\circ f)\,d\mu+\tfrac{1}{2}\int(u\circ g)\,d\mu$ (C.19)
$=\tfrac{1}{2}U(f)+\tfrac{1}{2}U(g)=U(f),$ (C.20)
and thus $\frac{1}{2}f+\frac{1}{2}g\succsim f$.
The next lemma is trivial.
Lemma 6 (Supermodularity-Submodularity Correspondence).
[…]
By [undefah], a capacity is submodular if and only if its associated Choquet integral is subadditive, i.e., for all bounded functions $h_{1},h_{2}$:
$\int(h_{1}+h_{2})\,d\mu\leq\int h_{1}\,d\mu+\int h_{2}\,d\mu.$ (C.21)
Now let $f\sim g$, i.e., $U(f)=U(g)$. By positive homogeneity and subadditivity of the Choquet integral:
$U\!\left(\tfrac{1}{2}f+\tfrac{1}{2}g\right)$ $=\int\left[\tfrac{1}{2}(u\circ f)+\tfrac{1}{2}(u\circ g)\right]d\mu\leq\tfrac{1}{2}\int(u\circ f)\,d\mu+\tfrac{1}{2}\int(u\circ g)\,d\mu$
$=\tfrac{1}{2}U(f)+\tfrac{1}{2}U(g)=U(f),$ (C.22)
and thus $\frac{1}{2}f+\frac{1}{2}g\precsim f$. ∎
Appendix D Proof of Proposition 1
@@
From the uniqueness properties of the biseparable preferences follows that $U^{1}=\theta U^{2}+\phi$, $u^{1}=\theta u^{2}+\phi$ and that the associated capacities are identical. What is left to show is the relation between the verifiable sets $\mathscr{V}^{1}$ and $\mathscr{V}^{2}$. Note that the sensitive events in the representation are exactly the sets $E$ for which $m^{\mu}(E)>0$ or for which there exists a cover of sets $E_{1},\ldots,E_{n}\subseteq E$ such that $\forall i:m^{\mu}(E_{i})>0$.
$cl_{\cup}\mathscr{V}^{1}\subseteq cl_{\cup}\mathscr{V}^{2}$: If $DV=cl_{\cup}(\mathscr{V}^{1})-cl_{\cup}(\mathscr{V}^{2})$ is nonempty, then by Lemma 1 (closure of sensitive events under unions) for every $\in DV, VV$ is sensitive in $U^{1}$. We show that $V$ cannot be sensitive in $U^{2}$: If $V$ is sensitive in $U^{2}$, then either $m^{\mu^{2}}(V)>0$ and thus $V\in\mathscr{V}^{2}$ or there exists a cover $V_{1},\ldots,V_{n}$ such that $\forall i:m^{\mu^{2}}(V_{i})>0$. If the latter is the case, then $V_{1},\ldots,V_{n}\in\mathscr{V}^{2}$ and thus $V\in cl_{\cup}(\mathscr{V}^{2})$. It follows that $V\in DV$ cannot be sensitive in $U^{2}$. But if $V$ is sensitive in $U^{1}$ but not in $U^{2}$, then the preferences cannot be identical.
$cl_{\cup}\mathscr{V}^{2}\subseteq cl_{\cup}\mathscr{V}^{1}$ follows by a symmetric argument.
@@
Appendix E Proof of Theorem 2